The following data relate to the operations of Shilow Company, a wholesale distributor of consumer goods: Current assets as oRequired: Using the preceding data: 1. Complete the schedule of expected cash collections. 2. Complete the merchandise purchaRequired 1 Required 2 Required 3 Required 4 Required 5 Complete the merchandise purchases budget and the schedule of expectedRequired 1 Required 2 Required 3 Required 5 Required 4 Complete the cash budget. (Cash deficiency, repayments and interest shRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare an absorption costing income statement for the quarter ended JRequired 5 Required 1 Required 2 Required 3 Required 4 Prepare a balance sheet as of June 30. Shilow Company Balance Sheet Ju
Answer #1
1) Shilow company
Schedule of Expected cash collections
April May June Quarter
Cash sales 51000 54000 69000 174000
credit sales 27,600 34000 36000 97,600
total collections 78600 88000 105000 271600
Accounts receivable = 115000*40%= 46000
2) Merchandise purchase budget
April May June Quarter
Budgeted cost of goods sold 63750 67500 86250 217500
Add Desired ending inventory 54000 69000 39,600 39,600
total needs 117750 136500 125850 257100
less beginning inventory 51,000 54,000 69,000 51,000
Required purchases 66,750 82,500 56,850 206,100
3) Schedule of Cash disbursements-Merchandise purhcase
April May June Quarter
March purchases 30,675 30,675
April purchases 33375 33,375 66750
May purchases 41250 41,250 82500
June purchases 28425 28425
total disbursements 64,050 74625 69675 208,350
4) Cash budget
April May June Quarter
Beginning cash balance 9,400 4,050 4,025 9,400
Add Cash collectiosn 78600 88000 105000 271600
total cas h available 88,000 92,050 109,025 281,000
less cash disbursements
for inventory 64,050 74625 69675 208,350
for expenses 19500 20400 24900 64800
for equipment 3,400 0 0 3,400
total cash disbursements 86,950 95025 94575 276,550
Excess(Deficiency)of cash 1,050 -2,975 14,450 4,450
Financing:
Borrowings 3,000 7,000 0 10,000
Repayments 0 -10,000 -10,000
interest 0 -230 -230
total financing 3,000 7,000 -10230 -230
Ending cash balance 4,050 4,025 4,220 4,220
interest = 3000*1%*3= 90
7000*1%*2= 140
230
5) income statement
Sales 290000
cost of goods sold
Beginning inventor 51,000
Add purchases 206,100
goods available for sale 257,100
ending inventory 39,600 217,500
Gross margin 72,500
Selling and administrative expense
commissions 34800
rent (4200*3) 12600
Depreciation (747*3) 2241
other expenses 17400 67041
net operating 5,459
interest expense 230
net income 5,229
Balance sheet
Assets
current assets
Cash 4,220
Accounts receivable 46,000
inventory 39,600
total current assets 89,820
Building And equipment ,net 100759
total Assets 190,579
liabilities And stockholder 's Equity
Accounts payable 28,425
total current assets 28,425
Stockholder's Equity
Capital stock 150,000
Retained earnings 12,154 162,154
total liabilites & stockholders Equity 190,579